Determine the atomic term symbols for the first excited state of the above atoms lithium to neon ( 3 < Z < 10).
Answer #1

The atomic term symbol is a description of angular momentum quantum numbers in a multi-electron atom. The general formula for the term symbols is 2s+1 LJ ;

S= total spin quantum number,

2S+1= Spin multiplicity

J = Total angular quantum numbers

L= Orbital quantum numbers

Term symbols for Lithium (Z=3):

The electronic configuration of Li is 1s22s1 and its term symbol is same as hydrogen that is 2S 1/2.

Term symbols for Beryllium (Z=4):

The electronic configuration of Be is 1s22s2 and its term symbol is same as helium that is 1S 0.

Term symbols for Boron (Z=5):

The electronic configuration of B is1s22s22p 1

Consider p orbital, whee l=1, L=1, s=1/2 hence, 2s+1=2, J=3/2 or 1/2. Therefore term symbol for boron is 2 P3/2 or 2P1/2.

Term symbols for Carbon (Z=6):

The electronic configuration of B is1s22s22p2. We consider only P electrons. therefore, l=1, hence ml = +1, 0, -1, L= 2,1,0 for D, P, and S term respectively. Here, two electrons spin parallell to each other, thus, S=0, therefore 2S+1 = 1, J=3,2, 1, 0. Hence the term symbol for carbon is 1D2, 3P2, 3P1, 3P0 and 1S0.   

Similarly for nitrogen the electronic configuration is 1s22s22p3 and its term symbol is 4S, 2P and 2D. The electronic configuration for oxygen is 1s22s22p4 . Here, consider that oxygen has 2P6 in the last orbit plus 2 positrons which forces them to take 2 electrons. Positrons differ in electron only in charge.Hence, the term arise form the np4 and np2 configuration is almost same and np5 is similar to np1.